OpenStudy (anonymous):

^3sqrt{sqrt(729x^4)}

OpenStudy (anonymous):

\[\sqrt{729x^2}=27x^2\] so \[\sqrt[3]{\sqrt{729x^4}}=\sqrt[3]{27x^2}\] \[=3\sqrt[3]{x^2}\]

OpenStudy (anonymous):

Another way of saying that is 3x^2/3